OpenStudy (anonymous):

How would you solve this equation: 1/2x^2=-32

OpenStudy (anonymous):

You would say that: \[\frac{1}{2x^2}=-32 \\ -\frac{1}{32}=2x^2 \\ -\frac{1}{64}=x^2 \\ x=\pm\sqrt{-\frac{1}{64}} \\ x=\pm\frac{i}{8}\] The solutions are imaginary which is understandable.

OpenStudy (anonymous):

After all, \[\frac{1}{2x^2}=-32\] can be rearranged to \[64x^2+1=0\] which is a parabola with no real solutions

OpenStudy (anonymous):

Thank you so much for helping me! However, my math book says that the answer is x= positive/negative 8i ... is this the same as your answer? Also, can you kindly tell me the second step in solving the equation in words

OpenStudy (anonymous):

The light bulb turned on! I know how to solve the answer now, thank you so very much for your help!

OpenStudy (anonymous):

No, that cant be the answer because if we substitute the answer, it is not correct: \[\frac{1}{2(\pm8i)^2}=-32 \\ \frac{1}{2(-64)}=-32 \\ -\frac{1}{128}=-32\] Of Course since there is a inequality, \(\pm8i\) cannot be a solution

OpenStudy (anonymous):

And no prob